OpenStudy (anonymous):

The sum of three mixed numbers is 20 4/15 Two of the numbers are 5 1/3 and 5 5/6 What is the third number? - Please teach me how to work this out! ><

OpenStudy (amistre64):

subtract the 2 given numbers from the total to find whats left over

OpenStudy (amistre64):

or we could simply set it up as such: \[5\frac{1}{3}+5\frac{5}{6}+N=20\frac{4}{15}\] and solve for N

OpenStudy (anonymous):

thank you :)!

OpenStudy (amistre64):

a mixed numeral is simply addition; if you have 5 and you add 1/3 of another to it; you get: 5 + 1/3 = 5' 1/3

OpenStudy (amistre64):

I would try it out like this: \[5\frac{1}{3}+5\frac{5}{6}+N=20\frac{4}{15}\] \[5+\frac{1}{3}+5+\frac{5}{6}+N=20+\frac{4}{15}\] \[5(3)+\frac{1(3)}{3}+5(3)+\frac{5(3)}{6}+(3)N=20(3)+\frac{4(3)}{15}\] \[15+1+15+\frac{5}{2}+3N=60+\frac{4}{5}\] \[31+\frac{5}{2}+3N=60+\frac{4}{5}\] \[31(2)+\frac{5(2)}{2}+3(2)N=60(2)+\frac{4(2)}{5}\] \[62+5+6N=120+\frac{8}{5}\] \[67+6N=120+\frac{8}{5}\] \[67+(-67)+6N=120+(-67)+\frac{8}{5}\] \[6N=53+\frac{8}{5}\] \[6(5)N=53(5)+\frac{8(5)}{5}\] \[30N=265+8\] \[30N=273\] \[\frac{30N}{30}=\frac{273}{30}\] \[N=\frac{273}{30}\] im sure there are steps you could do in your head, but it pretty much comes to this

OpenStudy (anonymous):

i got 9 3/30 but i don't think it's right.

OpenStudy (anonymous):

oh! i got it right when i put 9 1/10. THANKY YOU!

OpenStudy (amistre64):

youre welcome